OpenStudy (anonymous):

I asked a ? earlier and I still do not understand it could someone please help...Find the distance between these 2 points (-3,5) & (6,-2). Be sure that the answer is simplified and in radical form. Show all work.

hero (hero):

Plot the points and use pythagorean theorem or distance formula

OpenStudy (anonymous):

thats what the last person told me and apparently i am not doing something right

hero (hero):

You have graph paper right?

OpenStudy (anonymous):

no i am doing it all on computer

hero (hero):

Do you have a graphic plotting device on your computer?

OpenStudy (anonymous):

no we are not allowed to do it that way. I meant my assignment is on computer. I am having to do it all on paper and show all of my work

hero (hero):

you shouldn't do math without graph paper handy. Math 101

OpenStudy (anonymous):

ok I will remember that....i have the distance formula as sqrt((x2-x1)^2+(y2-y1)^2) but I do not know how to put everything into that formula or where to go from there

hero (hero):

Label the points and plug them in. It's easy

OpenStudy (anonymous):

ok let me see here give me a sec.

OpenStudy (anonymous):

when it comes to x1 and x2 does it matter which x # I put where

hero (hero):

yes it does

hero (hero):

label your points like this: (x1,y1) (x2, y2)

OpenStudy (anonymous):

ok so I would insert -3 in the x2 spot right

OpenStudy (anonymous):

ok now im confused

hero (hero):

okay, let me show you

hero (hero):

(x1,y1) = (-3,5) (x2,y2) = (6,-2) Is that clear enough?

OpenStudy (anonymous):

so when it says x2 or y2 that just means its place and not to the 2nd power

hero (hero):

yes, correct

OpenStudy (anonymous):

ok give me a sec

OpenStudy (anonymous):

Ok when I plug them all in I got sqrt((6-(-3))^2+(-2-(-5)^2=(3)^2+(3)^2=9+9=81

hero (hero):

You did something wrong. For that first one (6-(-3)^2 = (6+3)^2 = 9^2, not 3^2

hero (hero):

This is why graphing method is better

hero (hero):

But also 9+9 = 18 , not 81. You're supposed to add at the end not multiply

OpenStudy (anonymous):

Ok so the 2nd 1 would be (-2-(-5)^2=(7)^2=14

hero (hero):

You had the second one right already ....face palm :P

hero (hero):

(-2-(-5))^2 = (-2+5)^2 = (3)^2 = 9

OpenStudy (anonymous):

so my points would be (18,9)

hero (hero):

You're not trying to find anymore points. You're trying to find DISTANCE

OpenStudy (anonymous):

ok so what do i have to do from here

hero (hero):

Basically, when you calculated it the first time, you made a couple of adding mistakes. That's all. Try it again.

hero (hero):

And post your steps just like last time. Just be more careful

OpenStudy (anonymous):

what do you mean

hero (hero):

i meant for you to try to calculate it again. But try to add correctly this time.

OpenStudy (anonymous):

it would b 6 and not 9

hero (hero):

the answer is d = 3root(10). Try to figure out how I got that

OpenStudy (anonymous):

that would be 130